40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 574     Entry time: Mon Apr 11 18:20:02 2011
Author: tara 
Type: DailyProgress 
Category: BEAT 
Subject: beat setup is back 

The optics behind RCAV for beat measurement are re-installed, with the new periscope. I'm waiting for the temperature to settle.


The periscope behind RCAV was removed when we opened the chamber and took out the cavity. Now everything is back in place  

The new periscope is installed. The bottom mirror is still on the same mount we have used before(fig1), but the top part is removed, so it should be less sensitive to seismic compared to what it was before (fig2).


I removed the hose between the RCAV chamber and the turbo pump since the valve and the turbo were turned off. Then I closed the insulation box.

The yellow foam insulation on RCAV was fixed. I melted it a bit to make sure that no part of the beam is blocked by the insulation.



fig1: new periscope setup


 fig2) previous periscope

ELOG V3.1.3-